OpenStudy (anonymous):

Solve: y = - l x -3 l - 3

OpenStudy (karatechopper):

well the negative sign in front of those absolute values sign is -1. so if we remove the signs. u would multiply -1 by x and -3. so when u remove signs u get -x+3-3. which then if u simplify u get y=-x

OpenStudy (anonymous):

what does it mean "solve"?

OpenStudy (anonymous):

y is already solved for because it is by itself on one side of the equal sign. maybe it means "write as a piecewise function?"

OpenStudy (anonymous):

or maybe "solve" means "graph"?

OpenStudy (anonymous):

as a piecewise function you would get \[f(x) =- |x-3|-3 = \left\{\begin{array}{rcc} x & \text{if} & x \leq 3 \\ -x& \text{if} & x >3 \end{array} \right. \]

OpenStudy (karatechopper):

but what it was to just simplify?

OpenStudy (mr.math):

Good job Karate, you actually found one piece of the two pieces! :)

OpenStudy (mr.math):

in a very simple way!

OpenStudy (karatechopper):

uh ok